A train whistle ( f = 400Hz ) sounds higher or lower in frequency depending on whether it approaches or recedes. Calculate this difference for a train moving at a speed of 130km/h . Take the speed of sound in air to be 340m/s


1
Expert's answer
2021-02-23T14:01:31-0500

Answer

f=400Hz

V=130Km/hr=36m/s

"V_s=340m\/s"

When it goes towards to stationery observer

"f_a=f(\frac{V}{V-V_s}) =400(\frac{340}{340-36}) \\=447.4Hz"

Now when it's away from stationary obserser

"f_a'=f(\frac{V}{V+V_s}) =400(\frac{340}{340+36})\\ =361.7Hz"

Now frequency difference

"\Delta f=f_a-f_a'=447.4-361.7\\=85.7Hz"